The altitude of a certain right circular cone is the same as the radius of the base, and is measured as 5 inches with a possible error of 0.02 inch. Find approximately the percentage error in the calculated value of the volume.


1
Expert's answer
2021-01-20T02:46:28-0500

Let "h" be the altitude and "r" be the radius of the base of a right circular cone.

Given: "h=r=5" inches and "\Delta r=\Delta h=0.02" inches.

We know that the volume of the right circular is "V=\frac{1}{3}\pi r^2h"

Now "r=h\Rightarrow V=\frac{1}{3}\pi h^2h=\frac{1}{3}\pi h^3"

Require to find approximately the percentage error in the calculated value of the volume.

Taking logarithms on both sides of the equation "V=\frac{1}{3}\pi h^3", we get

"lnV=ln\left [ \frac{1}{3}\pi h^3 \right ]=ln\frac{\pi }{3}+ln\left ( h^3 \right )=ln\frac{\pi }{3}+3ln\left ( h \right )"

Taking differentials on both sides, we get

"\frac{\Delta V}{V}\approx 0+\frac{3\Delta h}{h}=\frac{3\Delta h}{h}"

Multiplying both sides by 100, we get

"\frac{\Delta V}{V}\times 100\approx \frac{3\Delta h}{h}\times 100"

Substituting "h=5" and "\Delta h=0.02" , we get

"\frac{\Delta V}{V}\times 100\approx \frac{3\left ( 0.02 \right )}{5}\times 100=1.2"

Therefore,

approximately the percentage error in the calculated value of the volume is "1.2"